# From Cavitation to Astrophysics: Explicit Solution of the Spherical Collapse Equation

### 12 Jan 2024 v2

@@ 3.2 Explicit solution $r(t)$
Figure 2: Exact solutions of the normalised collapse Equation (5) with initial conditions of Equation (6), evaluated on the normalised time interval $t\in[0,\tau(\gamma)]$ using the explicit Equation (26b). Coloured curves correspond to different values of $\gamma$, separated by $\Delta\gamma=1$. The dashed vertical line marks the collapse time $\sqrt{\pi/2}$, corresponding to the critical exponent $\gamma=-1$ that separates the divergent behaviour of $\dot{r}(\tau)$ from the convergent one.
3.3 Code examples
@@
The quantile function $Q(x;\alpha,\beta)$, which forms the heart of the explicit Equation (26a), is readily accessible in programming languages: qbeta in R, scipy.stats.beta.ppf in Python, InverseBetaRegularized in Mathematica, InvIncompleteBeta in the ALGLIB library for C++, C#, Java, Python, Delphi; etc. For reference, this section provides a few explicit code examples for plotting the collapse functions shown in Figure 2.
3.3.1 Gravitational spherical top-hat collapse
@@
3.3.2 Ideal spherical cavitation bubble collapse
Readers closer to cavitation dynamics than astrophysics may consider the example of a spherical cavity collapsing in an incompressible liquid of density $\rho$, without viscosity and surface tension. Assuming a constant driving pressure $\Delta p$, the bubble evolution is governed by Equation (10); hence the collapse motion $r(t)$ in natural units $R_{0}$ and $T_{0}=R_{0}\sqrt{\rho/\Delta p}$ is given by Equation (26a) with $\gamma=-4$. The corresponding R-script reads:
@@ 3.3.2 Ideal spherical cavitation bubble collapse
3.4 Noteworthy special solutions
For completeness, let us note that for select values of $\gamma$, Equation (26b) can be reduced to more compact explicit solutions:
$r(t)=$ $\sqrt{1-t^{2}}$ if $\gamma=-3$ (27a)
$r(t)=$ $2\displaystyle 1-t^{2}/$ 2 if $\gamma=0$ (27b)
$r(t)=$ $\cos(t)$ if $\gamma=1$ (27c)
$r(t)=$ ${\rm cn}\big{(}t,\tfrac{1}{2}\big{)}$ if $\gamma=3$, (27d)
where $\rm cn(x,y)$ is the Jacobi elliptic cosine function.
Equation (27d) is a special case of the undamped and unforced Duffing equation (Duffing, 1918), for which general solutions in terms of the Jacobi elliptic family have recently been derived by Salas et al. (2021).
For some other values of $\gamma$, parametric solutions, $t(\theta)$ and $r(\theta)$, have been found well before explicit solutions. Arguably, the most famous example is the case of $\gamma=-2$, describing, e.g., the spherical gravitational collapse (see Table 1). In this case (c.f., Lin et al., 1965),
Varying $\theta$ from 00 to $\pi$ generates the collapse phase, whereas varying it from $-\pi$ to 00 generates the symmetric growth.
More lengthy parametric solutions have also been presented for other non-trivial cases, such as empty spherical cavitation bubbles ($\gamma=-4$; c.f., Mancas & Rosu, 2016).
4 Discusion
The general solution $r(t)$ of Equations (5) and (6) given by Equation (26b) warrants a brief discussion. As a preliminary remark, the linear transformation (Equation 3) between the normalised form $r(t)$ and its dimensional analogue $R(T)$ make it straightforward to apply all properties of $r(t)$ to $R(t)$. Moreover, as mentioned in Section 1, any solution of Equations (1) and (2) is invariant under time reversal. This symmetry equally applies to dimensionless coordinates, i.e., $r(t)=r(-t)$, for all $t$, where a solution $r(t)$ exists. For efficiency, I will therefore limit the discussion in this section to the normalised form $r(t)$ and to positive times.
4.1 Behaviour at collapse point
As illustrated in Figure 2, the steepness of the function $r(t)$ monotonically increases as the time sweeps from $t=0$ ($r=1$) to the collapse time $t=\tau$ ($r=0$). Equation (8b) shows this immediately and reveals that in the limit $t\rightarrow\tau_{-}$ ($r\rightarrow 0_{+}$), the velocity $\dot{r}$ becomes
$\dot{r}(\tau)=\left\{\begin{array}[]{ll}-\sqrt{\frac{2}{1+\gamma}}&\text{if }\gamma>-1\\ -\infty&\text{if }\gamma\leq-1\end{array}\right..$ (29)
Some numerical values of $\dot{r}(\tau)$ are listed in Table 2. The divergent behaviour of $\dot{r}(\tau)$ for $\gamma\leq-1$, which can often be interpreted as a positively diverging kinetic energy ($\propto\dot{r}^{2}$), is mimicked by a negatively diverging potential $\phi(r)$. In fact, following Equation (12),
as visualised in Figure 1.
Equations (29) and (30) highlight the existence of two distinct regimes in the domain $\gamma\in\mathbb{R}$, separated by the critical value $\gamma=-1$. This value corresponds to the only singularity of Equation (26a), which can be properly removed by using Equation (26b). The dimensionless collapse time for $\gamma=-1$ is exactly $\sqrt{\pi/2}$ (Equation 24), the value shown by the dashed vertical line in Figure 2. Following Equation (29), all curves reaching $r=0$ to the left of this line do so vertically, whereas those to the right of the line come down at a finite slope.
The two regimes separated by $\gamma=-1$ often correspond to different classes of physical problems (c.f., Table 1), explaining why—to my knowledge—they have never been addressed simultaneously in previous literature.
The infinite velocity at the collapse point for $\gamma\leq-1$ is unphysical. Hence, all real-world systems modelled by Equation (1) must start deviating from this model near the collapse point. Secondary mechanisms, which might have been negligible for most of the collapse motion, suddenly become dominant, preventing the singularity. For example, in the case of collapsing cavitation bubbles, these mechanisms include shock waves (liquid compressibility), sonoluminescence, sonochemistry, vapour compression and micro-jetting (see overview by Obreschkow et al., 2013). In the case of matter collapsing by self-gravity, the mechanisms preventing the divergence could be asymmetries, smoothing the collapse point, and pressure forces, possibly enhanced by strong radiation and phase transitions, e.g., to a neutron star. If these mechanisms cannot prevent $\dot{R}(T)$ from approaching the speed of light, general relativistic effects take over, transforming the collapsing mass into a static black hole (Oppenheimer & Snyder, 1939). This is likely the fait of the cores of massive stars ($\gtrsim 25\,{\rm M}_{\odot}$ Mirabel, 2017) and possibly also much more massive ($\sim 10^{5}{\rm M}_{\odot}$) primordial gas clouds, collapsing directly into black holes (Loeb & Rasio, 1994).
4.2 Continuation past collapse point
For positive odd $\gamma$, however, $r(t)$ passes through $r(\tau)=0$, while inverting the sign of $\ddot{r}$. Symmetry considerations imply that $r(t)$ then describes an oscillating curve of wavelength $\lambda=4\tau$, made of reflections of the arc $r(t\in[0,\tau])$. This curve satisfies $r(0)=1$, $r(\tau)=0$, $r(2\tau)=-1$, $r(3\tau)=0$, $r(4\tau)=1$, and so forth. For $\gamma=1$, the oscillation is harmonic (Equation 27c), for all larger odd $\gamma$, it is anharmonic.
Non-real positive $\gamma\in\mathbb{R}_{+}$ would generally lead to complex solutions $r(t)\in\mathbb{C}$, if integrated past the collapse point. In fact, immediately past the collapse point, the acceleration has the complex argument $\arg(\ddot{r})=(\gamma+1)\pi$; hence the imaginary component of $\ddot{r}$ only vanishes for integer $\gamma$.
For negative $\gamma$ (integer or real), Equation (26b) diverges as $r\rightarrow 0$. This is a coordinate singularity, which can, in principle, be removed through regularisation techniques.
Figure 3: Dimensionless collapse time $\tau=T_{c}/T_{0}$ as a function of the control parameter $\gamma$. Plotted in semi-log coordinates, this function exhibits a non-trivial point symmetry about $(\gamma=-1,\tau=\sqrt{\pi/2})$, marked by the dashed lines. See Section 4.3.
4.3 $\gamma$$\tau$ Symmetry
Figure 3 shows the dimensionless collapse time $\tau$ as a function of the control parameter $\gamma$. With $\tau$ on a logarithmic axis, this function becomes an S-shape, which owes its symmetry to a well-known identity for Beta functions, $B(\alpha,\beta)B(\alpha+\beta,1-\beta)=\pi/(\alpha\sin(\pi\beta))$. Applied to Equation (21), this identity implies that any point $(\gamma,\tau(\gamma))$ can be mapped onto a corresponding point
which also lies on the function $\tau(\gamma)$. The only invariant point of this symmetry transformation, $(\gamma=-1,\tau=\sqrt{\pi/2})$, is marked by the dashed lines in Figure 3. This point coincides with the special case of Equation (24). It is the point, at which Equation (26a) becomes singular and where $\dot{r}(0)$ transitions from a finite to a diverging value (Section 4.1).
The symmetry Equation (31) does not have a straightforward known extension to general times $t\neq\tau$, say in the shape of $(\gamma,t,r)\mapsto(\gamma^{\prime},t^{\prime},r^{\prime})$. Hence, the two branches of solutions for $\gamma<-1$ (with diverging $\dot{r}$) and $\gamma>-1$ (with finite $\dot{r}$) cannot easily be expressed in terms of each other.
4.4 Polynomial approximations
$\tilde{r}(t)=$ $\!\left(1-x^{2}\right)^{p}$ if $\gamma\leq-1$ (32a)
$\tilde{r}(t)=$ $\!q(1\!-\!|x|)-(q-1)(1\!-\!|x|)^{\frac{q}{q-1}}$ if $\gamma>-1$, (32b)
Choosing the exponent $p$ equal to $p_{1}=2/(1-\gamma)$ (for $\gamma$ strictly below $-1$) ensures the correct asymptotic behaviour of $\dot{\tilde{r}}$ as a function of $\tilde{r}$ in the limit $t\rightarrow\tau$. This approximation was presented by Obreschkow et al. (2012) for $\gamma=-4$ (thus $p=2/5$). Equation (32a) can then be seen as the first term in a rapidly converging series, which tends towards the true solution if using the exact coefficients of Amore & Fernández (2013). Alternatively, setting $p$ equal to $p_{2}=\tau^{2}/2$, satisfies $\ddot{\tilde{r}}(0)=\ddot{r}(0)=-1$. Likewise, the parameter $q$ can be set to $q_{1}=\eta B(\eta,\tfrac{1}{2})$ (with $\eta$ defined in Equation 16) to ensure that $\dot{\tilde{r}}(\tau)=\dot{r}(\tau)$, or equal to $q_{2}=\tau^{2}/(\tau^{2}-1)$ to ensure $\ddot{\tilde{r}}(0)=\ddot{r}(0)=-1$.
As shown in Figure 4, both of the above choices for $p$ and $q$ provide reasonable approximations of the exact solution. In general, the best pick depends on the application. If $\gamma=-3$ (hence $p_{1}=p_{2}=\tfrac{1}{2}$) or $\gamma=0$ (hence $q_{1}=q_{2}=2$), the two options are identical and Equation (32b) is equal to the exact solutions of Equations (27a) and (27b), respectively.
A more quantitative discussion of the approximations is beyond the scope of this paper, but can be found in Obreschkow et al. (2012) and Amore & Fernández (2013) for $\gamma=-4$.
5 Conclusion
This paper investigated the differential Equation (1) with boundary conditions stated in Equation (2). One might call this the spherical collapse equation in reference to its most common applications (c.f., Table 1).
I have shown that this collapse equation exhibits a unified explicit solution, conveniently expressed in dimensionless coordinates $\{r=R/R_{0},t=T/T_{0}\}$. Elements of its derivation can be found elsewhere, spread over decades of literature across several unrelated fields. The contribution of this work is to unify previously known parametric, implicit and explicit solutions from hydrodynamics (c.f., Rayleigh-Plesset equation), hydrostatics (Lane-Emden equation) and astrophysics (top-hat spherical collapse model), all limited to different subsets of $\gamma$, into a single concise solution (Equation 26a), universally valid for all real $\gamma\neq-1$. For $\gamma=-1$ this solution is singular, but its limit $\gamma\rightarrow-1$ exists and is given in Equation (26b).
The existence of a general explicit solution is barely known, particularly in astrophysics and cosmology, where only parametric solutions like Equation (28) are commonly taught. I suspect this is because in the day when analytic argument was more mainstream, beta distributions would have been no easier to deal with than parametric solutions; and in the modern day, where beta distributions are readily accessible, the need for using them is less pressing, since numerical integration has become cheap.
The explicit solution $r(t)$ is not just more elegant and faster to evaluate than numerical integration, but it also offers: (i) a mathematical unification of seemingly unrelated physical applications (c.f., Table 1); (ii) interesting insights, for example pertaining to symmetry properties (Section 4.3); and (iii) a view on a hidden connection between the spherical collapse equation and probability theory, via the appearance of the beta distribution in Equation (26a). The explicit solution can provide exact benchmarks for testing numerical integration techniques and simulation codes, e.g., for hydrodynamic and gravitational simulations.
An interesting avenue for research in pure mathematics is an extension to complex-valued functions $r(t)$, where additional symmetries in the complex plane exist for $\gamma\in\mathbb{N}$.
Acknowledgements
I thank Dr. Martin Bruderer and Dr. Aaron Ludlow for insightful suggestions that have improved this work. I am also grateful to Dr. Roberto Iacono for bringing to my attention the convenience of using the regularised incomplete beta function in transitioning from Equation (22) to Equation (26a), bypassing a somewhat more cumbersome earlier method temporarily invoking the hypergeometric function. Last but not least, I extend my gratitude to Dr. Mohamed Farhat and Prof. Andrea Prosperetti for many past discussions, which were a source of inspiration in writing this paper. I am a recipient of an Australian Research Council Future Fellowship (FT190100083) funded by the Australian Government.
Data availability
No new data were generated or analysed in support of this research.

### 19 Mar 2024 v3

@@ III.2 Explicit solution $r(t)$
Figure 2: Exact solutions of the normalised collapse Equation (5) with initial conditions of Equation (6), evaluated on $t\in[0,\tau(\gamma)]$ using the explicit Equation (20b). Coloured curves correspond to different values of $\gamma$, separated by $\Delta\gamma=1$. The dashed vertical line marks the collapse time $\sqrt{\pi/2}$, corresponding to the critical exponent $\gamma=-1$ that separates the divergent behaviour of $\dot{r}(\tau)$ from the convergent one.
III.3 Noteworthy special solutions
For completeness, I note that for select values of $\gamma$, Equation (20a) can be reduced to well-known more compact solutions,
$r(t)=$ $\sqrt{1-t^{2}}$ if $\gamma=-3$ (21a)
$r(t)=$ $2\displaystyle 1-t^{2}/$ 2 if $\gamma=0$ (21b)
$r(t)=$ $\cos(t)$ if $\gamma=1$ (21c)
$r(t)=$ ${\rm cn}\big{(}t,\tfrac{1}{2}\big{)}$ if $\gamma=3$, (21d)
where ${\rm cn}(x,y)$ is the Jacobi elliptic cosine function.
Equation (21d) is a special case of the undamped and unforced Duffing equation [15], for which general solutions in terms of the Jacobi elliptic family have recently been derived [16].
For some other values of $\gamma$, parametric solutions have been found well before explicit solutions. A famous example is the case of $\gamma=-2$, describing, e.g., the spherical gravitational collapse (see Table 1). In this case [2],
Varying $\theta$ from 00 to $\pi$ generates the collapse phase, whereas varying it from $-\pi$ to 00 generates the symmetric growth phase. More lengthy parametric solutions have also been presented for other non-trivial cases, such as empty spherical cavitation bubbles ($\gamma=-4$) [17].
IV Discusion
The general solution of Equations (5) and (6) given by Equation (20b) warrants a brief discussion. As a preliminary remark, the linear transformation (Equation 3) between the normalised form $r(t)$ and its dimensional analogue $R(T)$ makes it straightforward to apply all properties of $r(t)$ to $R(t)$, and vice versa. For instance, as mentioned in Section I, any solution of Equations (1) and (2) is invariant under time reversal. This symmetry equally applies to dimensionless coordinates, i.e., $r(t)=r(-t)$, for all $t$, where a solution $r(t)$ exists. For efficiency, I will therefore limit the discussion in this section to the normalised form $r(t)$ and to positive times.
IV.1 Behaviour at collapse point
As illustrated in Figure 2, the steepness of the function $r(t)$ monotonically increases as the time sweeps from $t=0$ ($r=1$) to the collapse time $t=\tau$ ($r=0$). Equation (8b) shows this immediately and reveals that in the limit $t\rightarrow\tau_{-}$, the velocity $\dot{r}$ becomes
$\dot{r}(\tau)=\left\{\begin{array}[]{ll}-\sqrt{2/(1+\gamma)}&\text{if }\gamma>-1\\ -\infty&\text{if }\gamma\leq-1\end{array}\right..$ (23)
Some explicit values of $\dot{r}(\tau)$ are listed in Table 2.
The divergent behaviour of $\dot{r}(\tau)$ for $\gamma\leq-1$, which can often be interpreted as a positively diverging kinetic energy ($\propto\dot{r}^{2}$), is mimicked by a negatively diverging potential $\phi(r)$. In fact, following Equation (12),
as is apparent in Figure 1.
Equations (23) and (24) highlight the existence of two distinct regimes in the domain $\gamma\in\mathbb{R}$, separated by the critical value $\gamma=-1$. This value corresponds to the only singularity of Equation (20a), which is removed by Equation (20b). The dimensionless collapse time for $\gamma=-1$ is exactly $\sqrt{\pi/2}$ (Equation 18), the value shown by the dashed vertical line in Figure 2. Following Equation (23), all curves reaching $r=0$ to the left of this line do so vertically, whereas those to the right of the line come down at a finite slope.
The two regimes separated by $\gamma=-1$ correspond to different classes of physical problems (c.f., Table 1), explaining why—to my knowledge—they have never been addressed simultaneously in previous literature.
The infinite velocity at the collapse point for $\gamma\leq-1$ is unphysical, as all real-world systems modelled by Equation (1) start deviating from this model near the collapse point. Secondary mechanisms, which might have been negligible for most of the collapse motion, suddenly become dominant, preventing the singularity. For example, in the case of collapsing cavitation bubbles, these mechanisms include shock waves (liquid compressibility), sonoluminescence, sonochemistry, vapour compression and micro-jetting [18]. In the case of matter collapsing by self-gravity, the mechanisms preventing the divergence could be asymmetries, smoothing the collapse point, and pressure forces, possibly enhanced by strong radiation and phase transitions, e.g., to a neutron star. If these mechanisms cannot prevent $\dot{R}(T)$ from approaching the speed of light, general relativistic effects take over, transforming the collapsing mass into a static black hole [19]. This is likely the fate of the cores of massive stars ($\gtrsim 25\,{\rm M}_{\odot}$) [20] and possibly also primordial gas clouds ($\sim 10^{5}{\rm M}_{\odot}$), collapsing directly into black holes [21].
IV.2 Continuation past collapse point
For positive odd $\gamma$, however, $r(t)$ passes through $r(\tau)=0$, while inverting the sign of $\ddot{r}$. Symmetry considerations imply that $r(t)$ then describes an oscillating curve of wavelength $\lambda=4\tau$, made of reflections of the arc $r(t\in[0,\tau])$. This curve satisfies $r(0)=1$, $r(\tau)=0$, $r(2\tau)=-1$, $r(3\tau)=0$, $r(4\tau)=1$, and so forth. For $\gamma=1$, the oscillation is harmonic (Equation 21c), for all larger odd $\gamma$, it is anharmonic (e.g., Equation 21d).
Non-integer positive $\gamma$ generally lead to complex solutions $r(t)\in\mathbb{C}$, if integrated past the collapse point. In fact, immediately past this point, the acceleration has the complex argument $\arg(\ddot{r})=(\gamma+1)\pi$, implying that its imaginary part only vanishes for integer $\gamma$.
For negative $\gamma$, Equation (20b) diverges as $r\rightarrow 0$. This is a coordinate singularity, which can, in principle, be removed through regularisation techniques, such 2-body and $N$-body regularisation in astrophysical gravitational simulations [22].
Figure 3: Dimensionless collapse time $\tau=T_{c}/T_{0}$ as a function of the control parameter $\gamma$. Plotted in semi-log coordinates, this function exhibits a non-trivial point symmetry about $(\gamma=-1,\tau=\sqrt{\pi/2})$, marked by the dashed lines. See Section IV.3.
IV.3 Collapse time symmetry
Figure 3 shows the dimensionless collapse time $\tau$ as a function of the control parameter $\gamma$. With $\tau$ on a logarithmic axis, this function becomes a symmetric S-shape, which owes its symmetry to the beta function identity $B(\alpha,\beta)B(\alpha+\beta,1-\beta)=\pi/(\alpha\sin(\pi\beta))$. Applied to Equation (16), this identity implies that any point $(\gamma,\tau(\gamma))$ can be mapped onto a corresponding point
which also lies on the function $\tau(\gamma)$. This mapping is symmetric under the exchange of primed and non-primed variables. The only invariant point of this symmetry transformation, $\gamma=\gamma^{\prime}=-1$ and $\tau=\tau^{\prime}=\sqrt{\pi/2}$, is marked by the dashed lines in Figure 3. This point coincides with the special case of Equation (18), the singularity of Equation (20a) and the transition of $\dot{r}(\tau)$ from a finite to a diverging value (Section IV.1).
IV.4 General symmetry
A noteworthy property of Equation (5) is its formal invariance under a substitution $r\mapsto r^{\prime}:=r^{\delta}$, if $\delta=(1-\gamma)/2$, provable by invoking Equation (9). The restriction to collapse motions, implying $r,r^{\prime}\leq 1$ and $\ddot{r},\ddot{r}^{\prime}<0$, limits the applicability of this substitution to positive $\delta$, i.e., to $\gamma<1$. A few elementary steps show that this substitution corresponds to the transformation
$(\gamma,t,r)\mapsto\left(\gamma^{\prime}\!=\frac{\gamma+3}{\gamma-1},t^{\prime}\!=t\sqrt{\frac{1-\gamma}{2}},r^{\prime}\!=r^{\frac{1-\gamma}{2}}\right)\!.$ (26)
Like Equation (25), this mapping is symmetric under the exchange of primed and non-primed variables; and likewise this symmetry transformation exhibits a single invariant point at $\gamma=\gamma^{\prime}=-1$, where $(r^{\prime},t^{\prime})=(r,t)$.
The symmetry of Equation (25) can help identify new solutions based on existing ones, with the most trivial example being the derivation of Equation (21a) from Equation (21b), and vice versa.
When applying Equation (26) to dimensional quantities $R$ and $T$ instead of $r$ and $t$, the physical constant $k$ (Equation 1) must also be transformed, as $k^{\prime}=k\delta R_{0}^{1+\gamma}$. For example, Rayleigh’s Equation (10) can be written as
$\frac{{\rm d}^{2}Z}{{\rm d}t^{2}}=-\frac{5\Delta p}{2\rho}Z^{1/5},$ (27)
where $Z:=R^{5/2}$.
IV.5 Polynomial approximations
$\tilde{r}(t)=$ $\!\left(1-x^{2}\right)^{p}$ if $\gamma\leq-1$ (28a)
$\tilde{r}(t)=$ $\!q(1\!-\!|x|)\!-\!(q\!-\!1)(1\!-\!|x|)^{\frac{q}{q-1}}$ if $\gamma>-1$, (28b)
Let us first consider Equation (28a) (case $\gamma\leq-1$), where the exponent $p$ needs to be specified. Choosing $p$ equal to $p_{1}=2/(1-\gamma)$ (only for $\gamma$ strictly below $-1$) ensures the correct asymptotic behaviour of $\dot{\tilde{r}}$ as a function of $\tilde{r}$ in the limit $t\rightarrow\tau$. Earlier, I introduced this approximation specifically for $\gamma=-4$ (thus $p=2/5$) [7]. Equation (28a) can then be seen as the first term in a rapidly converging series, which tends toward the true solution if using exact analytical coefficients [8].
Alternatively, setting $p$ equal to $p_{2}=\tau^{2}/2$ (for all $\gamma\leq-1$), satisfies $\ddot{\tilde{r}}(0)=\ddot{r}(0)=-1$. For $\gamma=-2$, the case of the gravitational collapse, this choice is $p=(\pi/4)^{2}\approx 0.6169$. This is nearly identical to an analogous recent approximation [23], which reads $r(t)=(1-t^{2})^{p}$ with $p=1.8614/3\approx 0.6205$ determined via a fitting technique.
For Equation (28b) (case $\gamma>-1$), we have similar options for the parameter $q$. Setting it to $q_{1}=\eta B(\eta,\tfrac{1}{2})$ (with $\eta$ defined in Equation 13) ensures that $\dot{\tilde{r}}(\tau)=\dot{r}(\tau)$. Alternatively, setting it equal to $q_{2}=\tau^{2}/(\tau^{2}-1)$ ensures $\ddot{\tilde{r}}(0)=\ddot{r}(0)=-1$.
As shown in Figure 4, both of the above choices for $p$ and $q$ provide reasonable approximations of the exact solution. In general, the best pick depends on the application. If $\gamma=-3$ (hence $p_{1}=p_{2}=\tfrac{1}{2}$) or $\gamma=0$ (hence $q_{1}=q_{2}=2$), the two options are identical and Equations (28a) and (28b) become the exact solutions of Equations (21a) and (21b), respectively. Interestingly, these two solutions are related to each other via the symmetry transformation of Equation (26).
A more quantitative discussion of the approximations is beyond the scope of this paper, but can be found elsewhere for $\gamma=-4$ [7, 8].
V Conclusion
This paper investigated the differential Equation (1) with boundary conditions stated in Equation (2). I have called this the spherical collapse equation in reference to its most common applications (c.f., Table 1).
I have shown that this equation exhibits a unified explicit solution, invoking the beta distribution, a fundamental probability density function in statistics. Elements of the derivation can be found elsewhere, spread over decades of literature across several unrelated fields. The contribution of this work is to unify previously known parametric, implicit and explicit solutions from hydrodynamics (c.f., Rayleigh-Plesset equation), hydrostatics (Lane-Emden equation) and astrophysics (top-hat spherical collapse model), all limited to different subsets of $\gamma$, into a single concise solution (Equation 20a), universally valid for all real $\gamma\neq-1$. For $\gamma=-1$ this solution is singular, but its limit $\gamma\rightarrow-1$ exists, allowing the singularity to be removed via Equation (20b).
This general character ($\gamma\in\mathbb{R}$) of the explicit solution unifies seemingly unrelated physical applications. In this regard, this work provides context for recently found apparent analogies between cavitation bubbles and processes in astronomy and cosmology [24, 25, 26, 27].
The existence of a general explicit solution of Equation (1) is barely known, particularly in astrophysics and cosmology, where only parametric solutions like Equation (22) are commonly taught. Not only is the solution of Equation (20b) more elegant and faster to evaluate than numerical integration, but its explicit form can serve as a basis for further analytical manipulations and derivations to analyse particular problems. Illustrating this point, the theory of galaxies has benefited from Freeman’s analytic solution for the circular velocity of exponential disks [28]. This velocity can also be computed by direct numerical integration, but the explicit expression in terms of Bessel functions has greatly simplified further studies.
The explicit solution also offers insights pertaining to symmetry properties (Sections IV.3 and IV.4) and a connection between spherical collapse motions and probability theory. Finally, Equation (20b) can serve as an exact benchmark for testing numerical integration techniques and simulation codes, e.g., for hydrodynamic and gravitational simulations.
An avenue for research in pure mathematics is an extension to complex-valued functions $r(t)$, where additional symmetries in the complex plane exist for $\gamma\in\mathbb{N}$.
Acknowledgements.
I thank Dr. Martin Bruderer, Dr. Aaron Ludlow, Dr. Zachari Slepian, Dr. Dmitry Sinelshchikov and Prof. Stefani Mancas for insightful suggestions that have inspired and improved this work. I am also grateful to Dr. Roberto Iacono for bringing to my attention the convenience of using the regularised incomplete beta function in transitioning from Equation (15) to Equation (20a), bypassing an earlier method invoking the hypergeometric function. I also acknowledge an anonymous referee of an earlier draft, for pointing me toward the substitutions discussed in Section LABEL:sss:subsymmetry. Last but not least, I extend my gratitude to Dr. Mohamed Farhat and Prof. Andrea Prosperetti for many past discussions, which were a source of inspiration in writing this paper. I am a recipient of an Australian Research Council Future Fellowship (FT190100083) funded by the Australian Government.
Appendix A Analytical derivation
${\rm d}t^{2}=\frac{1+\gamma}{2(1-r^{1+\gamma})}{\rm d}r^{2},$ (29)
$\eta=\frac{1}{|1+\gamma|},$ (32)
$\alpha=\frac{1}{4}+\frac{3-\gamma}{4|1+\gamma|}.$ (35)
We have defined $\alpha$ in this way to make the right-hand integral of Equation (34) identical to the definition of the incomplete beta function, $B(x;\alpha,\beta)=\int_{0}^{x}y^{\alpha-1}(1-y)^{\beta-1}{\rm d}y$. Hence, Equation (34) solves to Equation (15).
Appendix B Code examples
@@
The quantile function $Q(x;\alpha,\beta)$, which forms the heart of the explicit Equation (20a), is readily accessible in most programming languages: qbeta in R, scipy.stats.beta.ppf in Python, BETA.INV in Excel, InverseBetaRegularized in Mathematica, InvIncompleteBeta in the ALGLIB library for C++, C#, Java, Python, Delphi; etc. For reference, this section provides a few explicit code examples for plotting the collapse functions shown in Figure 2.
B.1 Gravitational collapse
@@
B.2 Cavitation bubble collapse
Readers closer to cavitation dynamics than astrophysics may consider the example of a spherical cavity collapsing in an incompressible liquid of density $\rho$, without viscosity and surface tension. Assuming a constant driving pressure $\Delta p$, the bubble evolution is governed by Equation (10); hence the collapse motion $r(t)$ in natural units $R_{0}$ and $T_{0}=R_{0}\sqrt{\rho/\Delta p}$ is given by Equation (20a) with $\gamma=-4$. The corresponding R script reads: